OpenStudy (anonymous):

The age of the universe is about 5 X 10^17 seconds; the shortest light pulse produced in a laboratory (1990) lasted for only 6 X 10^-15second. Identify a physically meaningful time interval approximately halfway between these two on a logarithmic scale.

OpenStudy (anonymous):

log(10^2)=2 is 100 seconds, which is about how long I can lift weights. ;D

OpenStudy (jamesj):

Ok. To find the mid-point of the these two times on a log scale, first take their logs...

OpenStudy (jamesj):

Hence we have, using log base 10, the log of the two times is \[ x_1 = 17 + \log 5 , \ \ \ x_2 = \log 6 - 15 \] Now the mean of these two, the mid point, is just \[ \frac{x_1 + x_2}{2} = \frac{2 + \log 6 + \log 5}{2} \approx 1 + \log 6 \approx 1.5 \]

OpenStudy (jamesj):

So far so good?

OpenStudy (anonymous):

yea but book answer is 55 sec

OpenStudy (jamesj):

We haven't finished yet. Now to convert 1.5 back into a time unit we take the reverse operation; i.e., raise 10 to the power of 1.5 \[ 10^{1.5} = 31 \] So 31 seconds.

OpenStudy (jamesj):

Now if you did these calculations precisely, you'd find that the log-mean of the two numbers was closer to 1.739 and 10 raised to that power is 54.8 seconds

OpenStudy (jamesj):

Got it?

OpenStudy (jamesj):

It's good exercise in algebra to prove that this exactly equivalent to taking the geometric mean of these two numbers.

OpenStudy (anonymous):

why did you take a reverse operation when you find the time ?

OpenStudy (jamesj):

It's the definition of the "log-mean": - take the log of both numbers - find the mean (in the log scale) - reverse the log; i.e., take the exponent